# ProxyTransformation: Preshaping Point Cloud Manifold With Proxy Attention For 3D Visual Grounding

### v1

@@ 1 Introduction
In response, EmbodiedScan [45] was introduced, utilizing multi-view ego-centric RGB-D scans to enable models to process scenes directly from sparse, partial viewpoints. Previous methods[45, 68, 64] decouple the encoding of RGB images and depth-reconstructed point clouds from ego-centric perspectives to extract both semantic and geometric information, which are then projected into 3D space using intrinsic and extrinsic matrices to form a semantically enriched point cloud for bounding box regression. While effective, these methods overlook the noise, such as errors in depth on non-lambertian surfaces, captured by the depth sensors [20, 40, 21] that causes suboptimal performance. Consequently, critical geometric details of corresponding target objects are lost, potentially distorting their original manifold and further compromising the model’s grounding performance. Moreover, a substantial portion of the sampled points represents background regions, leading to a diminished density of foreground object points.
@@ 4.2 Analysis Experiments
Table 3: Ablation on Proxy Attention. Here, we compare the vanilla self-attention block with our Proxy Block. The FLOPs and parameters are computed over 3 transformer blocks with cluster features as input. We use a grid size of $12$ and a drop radio of $0.6$. The reported values are mAP for predictions greater than 25% IoU.
@@
Ablation on Proxy Attention. We validate the effectiveness and efficiency of proxy attention by comparing the vanilla attention block with the proxy block. The input cluster features are aligned across both blocks. As discussed in Sec. 3.2, the number of proxy tokens is significantly smaller than the input sequence length. This reduction in proxy tokens compresses redundancy within the attention module, enhancing computational efficiency, which aligns with our experiment results. As shown in Tab. 3, our approach achieves higher performance than vanilla attention, while reducing the computation overhead by 40.6%.
Ablation on Main Hyperparameters. Then, we ablate several hyperparameters in deformable point clustering and proxy transformation. As shown in Tab. 4, the size of the grid prior also influences performance by reducing instability during the early training stages, which can result from the randomness of the initial transformations. As shown in Tab. 5, we conduct several experiments to determine the optimal layers of our proxy attention block for efficient learning. Based on these findings, we selected the optimal hyperparameters for our experiments. Moreover, further ablation studies are given in the supplementary material.
4.3 Qualitative Results and Discussion
5 Conclusion

### v2

@@ 1 Introduction
In response, EmbodiedScan [46] was introduced, utilizing multi-view ego-centric RGB-D scans to enable models to process scenes directly from sparse, partial viewpoints. Previous methods[46, 70, 65, 66] decouple the encoding of RGB images and depth-reconstructed point clouds from ego-centric perspectives to extract both semantic and geometric information, which are then projected into 3D space using intrinsic and extrinsic matrices to form a semantically enriched point cloud for bounding box regression. While effective, these methods overlook the noise, such as errors in depth on non-lambertian surfaces, captured by the depth sensors [21, 41, 22] that causes suboptimal performance. Consequently, critical geometric details of corresponding target objects are lost, potentially distorting their original manifold and further compromising the model’s grounding performance. Moreover, a substantial portion of the sampled points represents background regions, leading to a diminished density of foreground object points.
@@ 4.2 Analysis Experiments
Table 3: Ablation on Proxy Attention. Here, we compare the vanilla self-attention block and cross-attention block with our Proxy Block. The FLOPs and parameters are computed over 3 transformer blocks with cluster features as input. We use a grid size of $12$ and a drop radio of $0.6$. The reported values are mAP for predictions greater than 25% IoU.
@@
Ablation on Proxy Attention. We validate the effectiveness and efficiency of proxy attention by comparing the vanilla attention block with the proxy block. The input cluster features are aligned across both blocks. As discussed in Sec. 3.2, the number of proxy tokens is significantly smaller than the input sequence length. This reduction in proxy tokens compresses redundancy within the attention module, enhancing computational efficiency, which aligns with our experiment results. As shown in Tab. 3, our approach achieves higher performance than vanilla attention, while reducing the computation overhead by 40.6%. And we have added an additional baseline that applies cross-attention as shown in Tab. 3, which demonstrate that our proposed Proxy Attention achieves a better trade-off between accuracy performance and computational efficiency.
Drop Radio Easy Hard Overall
0.5 40.55 34.17 40.03
0.6 41.66 34.38 41.08
0.7 40.67 32.70 40.02
0.8 39.65 31.02 38.95
0.9 36.40 27.02 35.64
Table 5: Ablation on Cluster Drop Radio $\beta$. A smaller $\beta$ indicates that a greater number of submanifolds are used for subsequent transformations. The reported values are mAP for predictions greater than 25% IoU.
Ablation on Main Hyperparameters. Then, we ablate several hyperparameters in deformable point clustering and proxy transformation. As shown in Tab. 4, the size of the grid prior also influences performance by reducing instability during the early training stages, which can result from the randomness of the initial transformations. As shown in Tab. 5, an excessively high drop ratio reduces model performance, highlighting the effectiveness of our submanifold transformations. A lower drop ratio retains more submanifolds for subsequent Proxy Transformation, enhancing results. However, experiments show that an overly low drop ratio also degrades performance, as point cloud enhancement should complement rather than overly alter the original structure. Moreover, further ablation studies are given in the supplementary material.
4.3 Qualitative Results and Discussion
5 Conclusion
Acknowledgement
The work is supported in part by the National Natural Science Foundation of China under Grant U24B2017, and Beijing Natural Science Foundation under Grant QY24257.
Appendix
A. Additional Experiments
Ablation on the number of Proxy Layers. As shown in Tab. 7, we conduct several experiments to determine the optimal layers of our proxy attention block for efficient learning. Based on these findings, we selected the optimal hyperparameters for our experiments. But we can see that even with just a single layer our model can achieve a SOTA performance. This demonstrates that the gain stems from leveraging multi-modal information rather than merely increasing model parameters. Thus, we can use a single-layer as final setup in practice, achieving SOTA results with minimal computational cost.
Ablation on $s$. As discussed when introducing deformable offsets, $s$ represents the maximum value of the offsets. Before generating the grid prior, we scale down the cuboid defined by the maximum and minimum coordinates of the point cloud based on $s$, ensuring that reference points, when adjusted by deformable offsets, do not move outside the boundaries of the point cloud. As shown in Tab. 6, we selected an optimal maximum value for $s$. An excessively large $s$ results in the reduced preset grid losing essential prior information, while an overly small $s$ prevents reference points from shifting toward more critical target regions, thereby reducing the flexibility of the model.
B. Ego-Centric 3D Visual Grounding
In real-world applications, intelligent agents interact with their surroundings without prior knowledge of the entire scene. Instead of relying on pre-reconstructed 3D point clouds or other scene-level priors commonly used in previous studies [50, 19], they primarily depend on ego-centric observations, such as multi-view RGB-D images.
Following the definition in [46], we formalize the ego-centric 3D visual grounding task as follows: Given a natural language query $L\in\mathbb{R}^{T}$, along with $V$ RGB-D image views $\{(I_{v},D_{v})\}_{v=1}^{V}$, where $I_{v}\in\mathbb{R}^{H\times W\times 3}$ denotes the RGB image and $D_{v}\in\mathbb{R}^{H\times W}$ represents the depth map for the $v$-th view, and their corresponding sensor intrinsics $\{(K^{I}_{v},K^{D}_{v})\}_{v=1}^{V}$ and extrinsics $\{(T^{I}_{v},T^{D}_{v})\}_{v=1}^{V}$, the goal is to predict a 9-degree-of-freedom (9DoF) bounding box $B=(x,y,z,l,w,h,\theta,\phi,\psi)$.
In this context, $(x,y,z)$ specify the 3D center coordinates of the target object, $(l,w,h)$ define its dimensions, and $(\theta,\phi,\psi)$ represent its orientation angles. The task is to determine $B$ such that it accurately localizes the object described by $L$ within the scene captured by $\{(I_{v},D_{v})\}_{v=1}^{V}$.
C. Details about Proxy Bias
As mentioned in the methodology, to compensate for the lack of positional information and the diversity of features, we propose a novel Proxy Bias:
$F=F_{0}+B^{\rm P},$ (16)
where $F\in\mathbb{R}^{N\times C}$ is the input of Proxy Block and $F_{0}\in\mathbb{R}^{N\times C}$ is our deformable cluster features. $B^{\rm P}\in\mathbb{R}^{N\times C}$ is our novel proxy bias.
Initially, we set three learnable parameters $B_{d}\in\mathbb{R}^{N\times D\times D}$, $B_{c}\in\mathbb{R}^{N\times 1\times S}$ and $B_{r}\in\mathbb{R}^{N\times S\times 1}$. Here, $C=S^{2}=D^{4}$. Therefore, our parameters are way less than directly setting $B^{\rm P}$ as a learnable parameter, thus improving our parameter efficiency.
We first interpolate $B_{d}$ into $B_{1}\in\mathbb{R}^{N\times S\times S}$, mapping the low-dimensional subspace into a higher-dimensional feature space to enhance feature diversity. Subsequently, we add $B_{c}$ and $B_{r}$ to obtain the final $B_{2}\in\mathbb{R}^{N\times S\times S}$, representing the linear union of two low-dimensional subspaces to form the final high-dimensional space, expressed as $V=V_{1}\cup V_{2}$. Finally, we get $B^{\rm P}=(B_{1}+B_{2}).{\rm reshape}(N,C)$, which can enrich the feature space with positional information and guide ProxyAttention to focus on diverse regions.
Max Easy Hard Overall
3 38.14 29.65 37.46
4 41.66 34.38 41.08
5 36.67 26.71 35.87
Table 6: Ablation on the Max Value of Offsets. The reported values are mAP for predictions greater than 25% IoU.